OpenStudy (anonymous):

Helpp! how many milliliters of a 10% hydrogen peroxide solution and how many milliliters of a 2% hydrogen peroxide solution should be mixed to obtain 500 mL of a 4% hydrogen peroxide solution?

OpenStudy (12man):

i will answer it

OpenStudy (ageta):

Equation: hydro + hydro = hydro 0.02x + 0.10(500-x) = 0.04*500 ------------------ 2x + 10*500 - 10x = 4*500 ---- -8x = -6*500 ---- x = (3/4)500 x = 3*125 = 375 ml (amt. of 2% solution needed) 500-x = 125 ml (amt. of 10% solution needed)

OpenStudy (12man):

yes @ageta is the answer

OpenStudy (ageta):

can i get a point

OpenStudy (anonymous):

what on earth? where did these answers come from?

OpenStudy (anonymous):

@AlysaMckee you have a clue as to how to do this?

OpenStudy (12man):

i dont have all day

OpenStudy (12man):

bye @ageta

OpenStudy (anonymous):

Not really @satellite73